# Inferring hyperuniformity from local structures via persistent homology

### v1

@@ 4 Inferring hyperuniformity from persistence diagrams
In this section, exploiting the information delivered by persistent homology introduced in section 3, we show that the HU character of a point cloud can be inferred from a characterization of its local arrangement. To this goal, we use standard machine learning (ML) methods [24], trained to encode the information contained in persistence diagrams. This is realized via selected features reconstructed from $h_{0}$ and $h_{1}$ diagram, presented below. We also outline the considered ML models and eventually report the results concerning the accuracy of the detection of hyperuniformity with the developed methods.
@@ 4.1 The features
Persistent homology characterized via persistence diagrams as introduced in section 3 cannot be readily used for models based on ML. First, ML methods typically work with information or features characterized by tens of values, while $h_{k}$ raw points are significantly more numerous: for $780$ points in the point clouds considered in this work, the persistence diagram would feature $\gtrsim 2000$ inputs ($779$ $h_{0}$ points and a number of $h_{1}$ points with the same order of magnitude); see the example in figure 3. Second, the number of $h_{1}$ points varies since the number of $1$-dimensional holes differs over different patterns: many ML models are not designed for receiving features characterized by a changing number of values. Therefore, generalized features that efficiently convert the data into a reduced and constant number of inputs are considered. We use three different ones. Two correspond to suitable features presented in reference [24], while we introduced an additional one accounting for quantifying distances between the entire $h_{1}$ persistence diagrams at once.
@@
The third feature is based on the Wasserstein distance that quantifies the distance between two point clouds $A$ and $B$, particularly useful in various fields such as statistics, ML, and optimal transport theory [34, 35]. We consider the $q$-th Wasserstein distance
@@
Adcock Algebraic VR: 69% Č: 69% VR: 92% Č: 90% VR: 92% Č: 69% VR: 92% Č: 69% VR: 92% Č: 69% VR: 91% Č: 89% VR: 91% Č: 89%
Kališnik Algebraic VR: 84% Č: 69% VR: 84% Č: 83% VR: 90% Č: 92% VR: 92% Č: 92% VR: 92% Č: 91% VR: 90% Č: 91% VR: 91% Č: 90%
@@ 4.2 The ML models
The training of the ML models was performed as follows: first, we extracted the features detailed in section 4.1 from the persistence diagrams of patterns with an ideal structure factor generated as described in section 2. Then, for each possible pair of features and ML models among the ones illustrated above, we have trained the model to detect hyperuniformity from the corresponding feature. To this goal, we labeled a point arrangement as HU or non-HU via a binary variable according to a condition on $\widetilde{H}$, i.e. the effective smaller value of the structure factor at the smallest sampled wavenumber; see equation (6). We consider a pattern HU if $\widetilde{H}<0.011$ approximating the nearly-HU condition $H\leq 0.01$ (as $\widetilde{H}$ typically overestimates $H$ for small $\alpha$, while approaching its value for large $\alpha$ and large systems only; see figure 1). The persistent homology leveraging Vietoris-Rips and the Čech complexes are considered, with independent training and testing phases. We used 5-fold cross-validation, i.e., we divided evenly and randomly the $1080$ patterns into 5 sets, and we trained the ML models $5$ times, training with $4$ of the sets and testing with the last one. This also enables us to compute a mean result as well as the standard deviation.
@@ 4.3 Results
The accuracy in the detection of the HU character in point clouds by varying features and ML models in the training is summarized in Table 1. The standard deviations range between $0.5\%$ and $3\%$ and are all reported in Table B1 in B. We can first observe that the hyperuniformity is overall well detectable from the local geometry of arrangements as described in persistence diagrams, up to an accuracy of $\gtrsim(96\pm 1.5)\%$. Such accuracy meets values typically considered very successful for ML algorithms combined with persistent homology [24]. These results thus unveil the existence of a very significant correlation between persistent homology, i.e., the local geometry of the arrangements in the patterns, and hyperuniformity. Thus, it shows that inverse criteria with respect to the systematic dependence of persistent homology on structure factor parametrization disclosed in reference [18] can be devised.
More precisely, we observe that the best results are obtained with Boosted Tree and Neural Networks ML models combined with the binning and the Wasserstein-distance-based features. This supports the usage of the latter, newly proposed here, for the analysis of HU arrangements and suggests its exploration for other kinds of patterns and applications as well. Both the different simplicial complexes considered to evaluate the persistent homology (Vietoris-Rips and Čech, see section 3) deliver similar performances over most of the considered models, except when considering Algebraic features. Being too dependent on the specific considered persistent homology, they should then be discarded for general applications of the method. It is worth mentioning that the linear SVM provides very good results with the binning feature up to $(94\pm 2)\%$ with $n\,$= 99\,99. Besides its simplicity, this points to a relatively robust linear separation in the feature space of HU and non-HU arrangements, which may be exploited for classification algorithms and generalizations based on classical (linear) regression methods.
@@ 5 Estimation of effective structure factor parameters from persistence diagram
$\mathcal{O}^{\prime}(\alpha,K)=\sum_{i=1}^{27}\left(D_{i}(\alpha,K)-[W]_{i}\right)^{2}$ (12)
performed here by exploiting the conjugate gradient repeated over different initial guesses for $\alpha,K$ to cope with the existence of local minima.
@@
For a broad range of parameters, the estimated $\alpha$ and $K$ values correspond well with the nominal values used to generate the patterns. Focusing first on the estimation of $\alpha$ for different nominal values of $K$, figure 6(a), we see that for intermediate nominal values $101<\alpha<$ 10, the prediction is very accurate with a narrow confidence interval. For the smallest sampled nominal $\alpha$, we observe that the estimation of $\alpha$ is accompanied by a broadening of the confidence interval. This points to less characteristic differences in the persistence diagrams of the corresponding patterns, expected due to a more significant contribution of randomness (see figure 1 for $\alpha\,$= 0.50.5\,0.50.5). Enough information is still retained to obtain estimation of $\alpha$ very close to nominal ones. Large nominal $\alpha$ values are also associated with a broadening of the confidence interval and, overall, a relatively accurate estimation of the nominal parameter when considering values within the confidence interval. This result may be ascribed to very similar topological arrangements realized for small variations of the parameters. Indeed, for large $K$ values (above $K_{\rm c}$, see section 2), nearly ordered patterns – namely triangular arrangements with defects – are obtained (see figure 1(a) for $K\,$= $\,56\pi$ and $\rm log(\alpha)\,$$>$ 11\,11). As such, minimal differences between corresponding persistence diagrams are realized when varying the structure factor parameters. This is also realized in the stealthy HU patterns obtained for smaller $K$ values.
@@
From these results, we may conclude that details set via global constraints in HU patterns can also be devised by looking at features of the local arrangements, with minor limitations for ordered and highly disordered systems only. Importantly, in the results of figure 6, the nominal values are almost always contained in the confidence interval. We remark that this interval, although arbitrary, points to the region with very similar arrangements. For example, in figure 6(b) for $\alpha\,$= 33\,33, the confidence interval has a width of $\approx$ $\,5\pi$ in $K$. Patterns generated with $\alpha\,$= 33\,33, $H\,$= $\,10^{-3}$, and the two different nominal values $K=40\pi$ and $K\,$= $\,45\pi$, together with their persistence diagrams, are shown in figure 7 and are clearly not easily distinguishable. Still, although featuring extremely similar geometrical arrangements and persistence diagrams, information concealed in the latter is enough to distinguish them and correctly estimate the parameter of their structure factor (as shown in figure 6(b)). Such a small but detectable difference obtained by varying $K$ by ±$\pm 5\pi$ is further used in the next section to assess the accuracy in generating a pattern approximating a given persistence diagram.
@@
Figure 7: Two point arrangements (first row), with nominal structure factor parameters reported above the panels, and (second row) their corresponding Čech persistence diagrams. Although they have a similar appearance, they are distinguished by the approach discussed in section 5; see figure 6(b), $\alpha=3.0$ panel.
@@ 6 Exploring the construction of patterns from (HU) persistence diagrams
We thus introduce a procedure to construct arrangements whose geometry approximates a given persistence diagram. For the sake of simplicity, we restrict the discussion to the Čech persistent homology, recalling that it usually contains more information than the Vietoris-Rips (see A), and consider the topological information encoded in $h_{1}$ diagrams. All the details are reported in C while we summarize here the main steps.
Figure 8: Illustration of the different basic constructions of triangles from $h_{1}$ points. $A$ is the point selected to apply the general construction algorithms (the marked point in the algorithms in C), $AB$ is a given edge with length $l_{\rm AB}$ and $p=(b,d)$ a given $h_{1}$ point. While (a) is the standard construction, (b) is applied when the edge $AB$ is too short for a correct application of the first one. If $AB$ is too large, (c) can be used if a $h_{1}$ point $\bar{p}$ satisfying the indicated conditions can be found. If not, (d) is applied. All these geometrical constructions are detailed in C. (e) An example of the recursive construction of triangles around a marked point, indicated in light blue. (f) Illustration of the closing procedure. By varying the position of $A^{\prime}$, we construct the AA’P (yellow) triangle to reproduce one of the targeted $h_{1}$ points, $p_{1}$. This implies the definition of the additional PA’B (green) triangle that is described by another $h_{1}$ point, $p_{2}$ (green squares in the persistence diagrams). While $p_{1}$ is in the targeted persistence diagram, $p_{2}$ is generally not. Different positions of $A^{\prime}$ and thus realizations of $p_{1}$ and $p_{2}$ are then inspected, and the one for which $p_{2}$ is the closest to one of the targeted $h_{1}$ points (red squares in the persistence diagrams) is selected.
Figure 9: Representative results demonstrating the generation of patterns with persistence diagrams that closely approximate the target ones. Three examples corresponding to the three rows are shown. (a) Original persistence diagrams –obtained here from ideal HU arrangements– to approximate. (b) Constructed point arrangements via the procedure outlined in section 6.1 from the persistence diagrams in panel (a). (c) Persistence diagrams of the point arrangements in panel (b), which approximate the ones in panel (a). The parameters of the patterns from which the persistence diagrams in panel (a) originate are from top to bottom: (i) $\alpha\,$= 1.01.0\,1.01.0, $H\,$= $\,5\cdot 10^{-4}$, $K\,$= 3434\,3434, (ii) $\alpha\,$= 3.03.0\,3.03.0, $H\,$= $\,5\cdot 10^{-4}$, $K\,$= 4545\,4545 and (iii) $\alpha\,$= 5.05.0\,5.05.0, $H\,$= $\,5\cdot 10^{-4}$ and $K\,$= 5656\,5656.
@@ 6.1 Construction algorithm
Using these methods, triangles are then constructed around a given point according to expected topological features; see, for instance, figure 8(e). The last iteration leading the point surrounded by triangles requires a closing procedure: different constructions are sampled, and the one associated with $h_{1}$ points closely resembling points in the targeted persistence diagram is selected. See figure 8(f) for an illustration. Once the point is surrounded by triangles, the same process is repeated on a peripheral point, i.e., a point not surrounded by triangles, where the peripheral point the closest to $(0,0)$ is chosen. The whole procedure is reported in detail in C: in Algorithm 1, the general method is described, while the construction of triangles and the closing procedures therein are summarized in the sub-Algorithms 2 and 3, respectively.
@@ 6.2 Results
We apply the algorithm outlined above to $18$ different persistence diagrams, all computed from different HU patterns with $H\,$= $\,5\cdot 10^{-4}$. Some examples of targeted persistence diagrams and generated patterns are shown in figure 9(a) and 9(b), respectively. The persistence diagram computed from these newly generated patterns, shown in 9(c), qualitatively well matches the original one, suggesting that the goal to approximate the targeted topological features has been reached. Some deviations between the persistence diagrams in 9(a) and 9(c) are still observed, such as “tails” along the diagonal. Note that without constraints on the region over which the points are generated and proceeding with local information only, the proposed construction method delivers a point arrangement that has a different global shape (a circle vs the square domain used to generate the ideal HU patterns as detailed in section 2). Consequently, and due to the locality of the construction, no boundary conditions are to be expected. Note that a consequence is that the structure factor of the constructed pattern inherently deviates from the one of the pattern from which the targeted persistence diagram is computed.
To quantitatively assess the accuracy of the algorithm, we compute the Wasserstein distance ($W_{1}$) of the original $h_{1}$ diagram (e.g. 9(a), with $\alpha_{0}$ and $K_{0}$ the parameters of the point pattern which generated it) to: i) the $h_{1}$ diagram of the reconstructed pattern (as, e.g., 9(c)), ii) the persistence $h_{1}$ diagram of a different ideal pattern generated with $\alpha\,$=$\,\alpha_{0}$, $K\,$=$\,K_{0}$, iii) the persistence $h_{1}$ diagrams of patterns generated with $\alpha\,$=$\,\alpha_{0}$, $K\,$=$\,K_{0}\pm 5$ (mimicking small variations as in figure 7). The results summarized in figure 10 show that the algorithm performs very well quantitatively. Persistence $h_{1}$ diagrams of the original patterns have a $W_{1}$ distance from the diagrams of another ideally generated pattern (green) similar to the constructed one (red). The distances from the persistence $h_{1}$ diagrams for patterns with $\alpha\,$=$\,\alpha_{0}$, $K\,$=$\,K_{0}\pm 5$ (implying only tiny geometrical changes as explained in section 5) always largely exceed these values. This is obtained for all the analyzed cases. Still, consistent with the results of the previous section, the algorithm is found to perform very well for disordered HU patterns, while it shows some larger deviation from ideality for large $\alpha$, particularly when leading to ordered systems at large $K$ (see the $K\,$= $\,56\pi$ panel in figure 10).

### v2

@@ 4 Inferring hyperuniformity from persistence diagrams
In this section, we show that the HU character of a point pattern can be inferred from a characterization of its local arrangements via persistent homology (introduced in section 3). To this goal, we use standard machine learning (ML) methods [31], trained to encode the information contained in persistence diagrams. This is realized via selected features reconstructed from $h_{0}$ and $h_{1}$ diagram, presented below. We also outline the considered ML models and eventually report the results concerning the accuracy of the detection of hyperuniformity with the developed methods.
@@ 4.1 The features
Persistent homology characterized via persistence diagrams as introduced in section 3 cannot be readily used for models based on ML. First, ML methods typically work with information or features characterized by tens of values, while raw $h_{k}$ points are significantly more numerous: for $780$ points in the point patterns considered in this work, the persistence diagram would feature $\gtrsim 2000$ inputs ($779$ $h_{0}$ points and a number of $h_{1}$ points with the same order of magnitude); see the example in figure 3. Second, the number of $h_{1}$ points varies since the number of $1$-dimensional holes differs over different patterns: many ML models are not designed for receiving features characterized by a changing number of values. Therefore, generalized features that efficiently convert the data into a reduced and constant number of inputs are considered. We use three different ones. Two correspond to suitable features presented in reference [31], while we introduced an additional one accounting for quantifying distances between the entire $h_{1}$ persistence diagrams at once.
@@
The third feature is based on the Wasserstein distance that quantifies the distance between two sets of points $A$ and $B$, particularly useful in various fields such as statistics, ML, and optimal transport theory [45, 46]. We consider the $q$-th Wasserstein distance
@@
Adcock Algebraic VR: 72% Č: 72% VR: 92% Č: 90% VR: 92% Č: 72% VR: 92% Č: 72% VR: 92% Č: 72% VR: 91% Č: 89% VR: 91% Č: 89%
Kališnik Algebraic VR: 84% Č: 72% VR: 84% Č: 83% VR: 90% Č: 92% VR: 92% Č: 92% VR: 92% Č: 91% VR: 90% Č: 91% VR: 91% Č: 90%
@@ 4.2 The ML models
The training of the ML models was performed as follows: first, we extracted the features detailed in section 4.1 from the persistence diagrams of patterns with an ideal structure factor generated as described in section 2. Then, for each possible pair of features and ML models among the ones illustrated above, we have trained the model to detect hyperuniformity from the corresponding feature. To this goal, we labeled a point pattern as HU or non-HU via a binary variable according to a condition on $\widetilde{H}$, i.e. the effective smaller value of the structure factor at the smallest sampled wavenumber; see equation (6). We consider a pattern HU if $\widetilde{H}<0.011$ approximating the nearly-HU condition $H\leq 0.01$ (as $\widetilde{H}$ typically overestimates $H$ for small $\alpha$, while approaching its value for large $\alpha$ and large systems only; see figure 1). The persistent homology leveraging Vietoris-Rips and the Čech complexes are considered, with independent training and testing phases. We used 5-fold cross-validation, i.e., we divided evenly and randomly the $1080$ patterns into 5 sets, and we trained the ML models $5$ times, training with $4$ of the sets and testing with the last one. This also enables us to compute a mean result as well as the standard deviation.
@@ 4.3 Results
The accuracy in the detection of the HU character in point patterns by varying features and ML models in the training is summarized in Table 1. The standard deviations range between $0.5\%$ and $3\%$ and are all reported in Table B1 in B. We can first observe that the hyperuniformity is overall well detectable from the local geometry of arrangements as described in persistence diagrams, up to an accuracy of $(96\pm 1.5)\%$. Such accuracy meets values typically considered very successful for ML algorithms combined with persistent homology [31]. These results thus unveil the existence of a very significant correlation between persistent homology, i.e., the local geometry of the arrangements in the patterns, and hyperuniformity. Thus, it shows that inverse criteria with respect to the systematic dependence of persistent homology on structure factor parametrization disclosed in reference [23] can be devised.
More precisely, we observe that the best results are obtained with Boosted Tree and Neural Networks ML models combined with the binning and the Wasserstein-distance-based features. This supports the usage of the latter, newly proposed here, for the analysis of HU arrangements and suggests its exploration for other kinds of patterns and applications as well. Both the different simplicial complexes considered to evaluate the persistent homology (Vietoris-Rips and Čech, see section 3) deliver similar performances over most of the considered models, except in the case of Algebraic features. Being too dependent on the specific considered persistent homology, they should then be discarded for general applications of the method. It is worth mentioning that the linear SVM provides very good results with the binning feature up to $(94\pm 2)\%$ with $n\,$= 99\,99. Besides its simplicity, this points to a relatively robust linear separation in the feature space of HU and non-HU arrangements, which may be exploited for classification algorithms and generalizations based on classical (linear) regression methods.
Although we focused here on a classification problem, features based on persistence diagrams and suitable ML can be applied to regression problems as well, provided suitable data sets are available and/or can be generated. In C, we show that such a framework allows for estimating the value of the parameter $H$ directly, confirming the possibility of inferring detailed characteristics of hyperuniformity from persistence diagrams.
@@ 5 Estimation of effective structure factor parameters from persistence diagram
$\mathcal{O}^{\prime}(\alpha,K)=\sum_{i=1}^{27}\left(D_{i}(\alpha,K)-[W]_{i}\right)^{2},$ (12)
performed here by exploiting the conjugate gradient method, repeated over different initial guesses for $\alpha,K$ to cope with the existence of local minima.
@@
For a broad range of parameters, the estimated $\alpha$ and $K$ values correspond well with the nominal values used to generate the patterns. Focusing first on the estimation of $\alpha$ for different nominal values of $K$, figure 6(a), we see that for intermediate nominal values $101<\alpha<$ 10, the prediction is very accurate with a narrow confidence interval. For the smallest sampled nominal $\alpha$, we observe that the estimation of $\alpha$ is accompanied by a broadening of the confidence interval. This points to less characteristic differences in the persistence diagrams of the corresponding patterns, expected due to a more significant contribution of randomness (see figure 1 for $\alpha\,$= 0.50.5\,0.50.5). Enough information is still retained to obtain estimation of $\alpha$ very close to nominal ones. Large nominal $\alpha$ values are also associated with a broadening of the confidence interval and, overall, a relatively accurate estimation of the nominal parameter when considering values within the confidence interval. This result may be ascribed to very similar topological arrangements realized for small variations of the parameters. Indeed, for large $K$ values (above $K_{\rm c}$, see section 2), nearly ordered patterns – namely triangular arrangements with defects – are obtained (see figure 1(a) for $K\,$= $\,51\pi$ and $\rm\alpha=100$). As such, minimal differences between corresponding persistence diagrams are realized when varying the structure factor parameters. This is also realized in the stealthy HU patterns obtained for smaller $K$ values.
@@
From these results, we may conclude that details set via global constraints in HU patterns can also be devised by looking at features of the local arrangements, with minor limitations for ordered and highly disordered systems only. Importantly, in the results of figure 6, the nominal values are almost always contained in the confidence interval. We remark that this interval, although arbitrary, points to the region with very similar arrangements. For example, in figure 6(b) for $\alpha\,$= 33\,33, the confidence interval has a width of $\approx$ $\,5\pi$ in $K$. Patterns generated with $\alpha\,$= 33\,33, $H\,$= $\,10^{-3}$, and the two different nominal values $K=40\pi$ and $K\,$= $\,45\pi$, together with their persistence diagrams, are shown in figure 7 and are clearly not easily distinguishable. Still, although featuring extremely similar geometrical arrangements and persistence diagrams, information concealed in the latter is enough to distinguish them and correctly estimate the parameter of their structure factor (as shown in figure 6(b)). Such a small but detectable difference obtained by varying $K$ by ±$\pm 5\pi$ is further used in the following to assess the accuracy in generating a pattern approximating a given persistence diagram.
@@
Figure 7: Two point patterns (first row), with nominal structure factor parameters reported above the panels, and (second row) their corresponding Čech persistence diagrams. Although they have a similar appearance, they are distinguished by the approach discussed in section 5; see figure 6(b), $\alpha=3.0$ panel.
6 Detecting boundaries between disordered arrangements
A scenario that can be handled efficiently by persistent homology is the coexistence of different point patterns. It resembles a situation very common in ordered systems like (poly)crystals, where domains with a different order or orientations (grains) are separated by defected regions (grain boundaries) [19]. Inhomogeneities emerge at the boundaries of disordered systems, which we show below are captured by the information contained in persistence diagrams.
Figure 8: Four-grains pattern. (a) construction method: four different homogeneous patterns sharing the same structure factor (here $\alpha=3.0$, $H=10^{-3}$ and $K=51\pi$) are stitched together, and a central square containing $780$ points (dark grey) is selected. (b) Resulting four-grains pattern (dark grey in panel a).
A new arrangement exhibiting boundaries between point patterns featuring ideal HU characters is generated. Four different realizations of ideal HU patterns (780 points) with the same structure factor are juxtaposed into a large square pattern with $4\times 780$ points. To have a comparable persistence diagram between this new “four-grain” setting and the ideal HU patterns (“one-grain”), a square region centered in the middle and containing $780$ points is selected. This construction is illustrated in figure 8. We focus here on patterns featuring a marked HU character with $\alpha\geq 3$, $H=10^{-3}$ and $K\geq 45\pi$.
To quantify the difference between the persistence diagrams of four- and one-grain patterns, we use the Wasserstein distance as introduced in equation (10). More precisely, we evaluated this distance 12 times with different four-grain and one-grain realizations, obtaining a mean value as well as a standard deviation. These results can then be compared to the typical distance between two one-grain patterns (as in previous sections) extracted from 6 different evaluations. The obtained standard deviation intervals are summarized in figure 9. They clearly show that the presence of boundaries is encoded in the persistence diagrams by a significant increase in their (Wasserstein) distance from those of ideal, one-grain patterns. As an illustration, the case with $\alpha=3.0$ and $K=51\pi$ presented in figure 8 induces a difference of around $20$% of the distances. In other words, the presence of boundaries between “grains” is effectively captured by the persistent homology.
This result further confirms the suitability of the considered method for the study of local geometry beyond homogeneous arrangements. Moreover, by focusing on subsets of inhomogeneous patterns, methods can be envisaged to detect where the boundary regions are located, if any. Cases involving overlap of patterns of empty space between patterns would lead to even more significant features in the persistence diagrams and can be treated similarly to the case explicitly discussed in this section.
Figure 9: Wasserstein distance between the persistence diagrams of one-grain from both other one- and four-grain patterns (blue-green and red regions, respectively). Colored regions are centered on the mean Wasserstein distance with a width equal to the standard deviation over several independent evaluations.
@@ 7 Exploring the construction of patterns from (HU) persistence diagrams
We thus introduce a procedure to construct arrangements whose geometry approximates a given persistence diagram. For the sake of simplicity, we restrict the discussion to the Čech persistent homology, recalling that it usually contains more information than the Vietoris-Rips (see A), and consider the topological information encoded in $h_{1}$ diagrams. All the details are reported in D while we summarize here the main steps.
Figure 10: Illustration of the different basic constructions of triangles from $h_{1}$ points. $A$ is the point selected to apply the general construction algorithms (the marked point in the algorithms in D), $AB$ is a given edge with length $l_{\rm AB}$ and $p=(b,d)$ a given $h_{1}$ point. While (a) is the standard construction, (b) is applied when the edge $AB$ is too short for a correct application of the first one. If $AB$ is too large, (c) can be used if a $h_{1}$ point $\bar{p}$ satisfying the indicated conditions can be found. If not, (d) is applied. All these geometrical constructions are detailed in D. (e) An example of the recursive construction of triangles around a marked point, indicated in light blue. (f) Illustration of the closing procedure. By varying the position of $A^{\prime}$, we construct the AA’P (yellow) triangle to reproduce one of the targeted $h_{1}$ points, $p_{1}$. This implies the definition of the additional PA’B (green) triangle that is described by another $h_{1}$ point, $p_{2}$ (green squares in the persistence diagrams). While $p_{1}$ is in the targeted persistence diagram, $p_{2}$ is generally not. Different positions of $A^{\prime}$ and thus realizations of $p_{1}$ and $p_{2}$ are then inspected, and the one for which $p_{2}$ is the closest to one of the targeted $h_{1}$ points (red squares in the persistence diagrams) is selected.
Figure 11: Representative results demonstrating the generation of patterns with persistence diagrams that closely approximate the target ones. Three examples corresponding to the three rows are shown. (a) Original persistence diagrams –obtained here from ideal HU arrangements– to approximate. (b) Constructed point patterns via the procedure outlined in section 7.1 from the persistence diagrams in panel (a). (c) Persistence diagrams of the point patterns in panel (b), which approximate the ones in panel (a). The parameters of the patterns from which the persistence diagrams in panel (a) originate are from top to bottom: (i) $\alpha\,$= 1.01.0\,1.01.0, $H\,$= $\,5\cdot 10^{-4}$, $K\,$= 3434\,3434, (ii) $\alpha\,$= 3.03.0\,3.03.0, $H\,$= $\,5\cdot 10^{-4}$, $K\,$= 4545\,4545 and (iii) $\alpha\,$= 5.05.0\,5.05.0, $H\,$= $\,5\cdot 10^{-4}$ and $K\,$= 5656\,5656.
@@ 7.1 Construction algorithm
Using these methods, triangles are then constructed around a given point according to expected topological features; see, for instance, figure 10(e). The last iteration leading the point surrounded by triangles requires a closing procedure: different constructions are sampled, and the one associated with $h_{1}$ points closely resembling points in the targeted persistence diagram is selected. See figure 10(f) for an illustration. Once the point is surrounded by triangles, the same process is repeated on a peripheral point, i.e., a point not surrounded by triangles, where the peripheral point the closest to $(0,0)$ is chosen. The whole procedure is reported in detail in D: in Algorithm 1, the general method is described, while the construction of triangles and the closing procedures therein are summarized in the sub-Algorithms 2 and 3, respectively.
@@ 7.2 Results
We apply the algorithm outlined above to $18$ different persistence diagrams, all computed from different HU patterns with $H\,$= $\,5\cdot 10^{-4}$. Some examples of targeted persistence diagrams and generated patterns are shown in figure 11(a) and 11(b), respectively. The persistence diagram computed from these newly generated patterns, shown in 11(c), qualitatively well matches the original one, suggesting that the goal to approximate the targeted topological features has been reached. Some deviations between the persistence diagrams in 11(a) and 11(c) are still observed, such as “tails” along the diagonal. Note that without constraints on the region over which the points are generated and proceeding with local information only, the proposed construction method delivers a point pattern that has a different global shape (a circle vs the square domain used to generate the ideal HU patterns as detailed in section 2). Consequently, and due to the locality of the construction, no boundary conditions are to be expected. Note that a consequence is that the structure factor of the constructed pattern inherently deviates from the one of the pattern from which the targeted persistence diagram is computed.
To quantitatively assess the accuracy of the algorithm, we compute the Wasserstein distance ($W_{1}$) of the original $h_{1}$ diagram (e.g. 11(a), with $\alpha_{0}$ and $K_{0}$ the parameters of the point pattern which generated it) to: i) the $h_{1}$ diagram of the reconstructed pattern (as, e.g., 11(c)), ii) the persistence $h_{1}$ diagram of a different ideal pattern generated with $\alpha\,$=$\,\alpha_{0}$, $K\,$=$\,K_{0}$, iii) the persistence $h_{1}$ diagrams of patterns generated with $\alpha\,$=$\,\alpha_{0}$, $K\,$=$\,K_{0}\pm 5$ (mimicking small variations as in figure 7). The results summarized in figure 12 show that the algorithm performs very well quantitatively. Persistence $h_{1}$ diagrams of the original patterns have a $W_{1}$ distance from the diagrams of another ideally generated pattern (green) similar to the constructed one (red). The distances from the persistence $h_{1}$ diagrams for patterns with $\alpha\,$=$\,\alpha_{0}$, $K\,$=$\,K_{0}\pm 5$ (implying only tiny geometrical changes as explained in section 5) always largely exceed these values. This is obtained for all the analyzed cases. Still, consistent with the results of the previous section, the algorithm is found to perform very well for disordered HU patterns, while it shows some larger deviation from ideality for large $\alpha$, particularly when leading to ordered systems at large $K$ (see the $K\,$= $\,56\pi$ panel in figure 12).